We roll a die until the sum of points obtained exceeds 350. Approximate the probability that we will roll more than 120 times.
Answer #1

here for a single roll probability of an outcome is 1/6

x P(x) xP(x) x2P(x)
1 1/6 0.167 0.167
2 1/6 0.333 0.667
3 1/6 0.500 1.500
4 1/6 0.667 2.667
5 1/6 0.833 4.167
6 1/6 1.000 6.000
total 3.500 15.167
E(x) =μ= ΣxP(x) = 3.5000
E(x2) = Σx2P(x) = 15.1667
Var(x)=σ2 = E(x2)-(E(x))2= 2.917
std deviation=         σ= √σ2 = 1.7078

hence for 120 rolls; expected sum of numbers =120*3.5=420

and std deviation =1.7078*sqrt(120)=18.71

hence probability that we will roll more than 120 times untill sum exceed 350

=P(in 120 rolls sum reached at most 350)=P(X<=350)=P(Z<(350.5-420)/18.71)=P(Z<-3.71)

=0.0001